OpenStudy (anonymous):

Solve x4 – 10x2 + 9 = 0. x = ±1 or x = ±3 x = ±3 or x = ±i x = ±1 or x = mc074-1.jpg x = ±3 or x = mc074-2.jpg

OpenStudy (anonymous):

(x2 -9)(x2 -1) = 0

OpenStudy (anonymous):

set your factors to zero and you have your answer

OpenStudy (anonymous):

so i.e. would be 3 or 1?

OpenStudy (anonymous):

(x2 -9) = 0 , (x2 -1) = 0 Solve for x on both